OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL! PLEASE HELP!) By making a change of variables and letting a = 1/x and b = 1/y, the following system can be converted into a system of linear equations in terms of a and b. Solve the new system and use your solution to find the solution to the original system of equations: 2/x - 3/y = 11/3 3/x + 2/y = -1

OpenStudy (anonymous):

So, plug in a and b where 1/x and 1/y are, then solve. Where are you having trouble?

OpenStudy (anonymous):

I'm getting the equations 6a - 9b - 33 and 3a + 2b = 11 when I do so, is that correct?

OpenStudy (anonymous):

I'm not exactly sure how you got that... \[2*\frac{ 1 }{ x }- 3*\frac{1}{y}=\frac{11}{3}\] and \[3*\frac{1}{x}+2*\frac{1}{y}=-1\] right? So just replace 1/x and 1/y with a and b respectively.

OpenStudy (anonymous):

I'll try that, thank you!

OpenStudy (anonymous):

I'm getting x = 3 and y = -1?

OpenStudy (anonymous):

That's what I'm getting too,

OpenStudy (anonymous):

Wait I figured it out! Thank you so much for your help though!